A lunar landing module is descending to the moon`s surface at a steady velocity of 10m/s at a height of 120m , a small object falls from its landing gear . Taking the moon`s gravitational acceleration as 1.6m/s at what speed in m/s does the object strike the moon ?

Grade:

1 Answers

chaitnyakishore
10 years ago
Given initial velocity of the object is `v2`=10m/s, height at which the object falls h=120m, Let the final velocity of the object when it strikes the surface be `v1` and acceleration due to gravity on moon(g`)=1.6m/s² By kinematic equations: (v2)²-(v1)²=2as Here acceleration is replaced by acceleration due to gravity of moon and distance is replaced by height and since the object is heading towards moon the g is taken negative. Therefore 10²-(v1)²=-2gh 100-(v1)²=-2*1.6*120 (v1)²=100+384 (v1)²=484 v1=(v484) v1=22m/s Therefore velocity of the object when it strikes the surface of the moon is `v1`=22m/s